OpenStudy (anonymous):

a secretary types 4160 words an hpur and 20 minutes. At the same rate how many words can be typed in an 8 hour day assuming no breaks? _______ words Pleaseee Help

OpenStudy (anonymous):

8*60=480 480/80=6 6*4160=24960

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

uw

OpenStudy (anonymous):

can you help me with a few more?